Please answer all steps/questions for this problem, round if needed. Please and thank you! :) An engineer is going to redesign an ejection seat for an airplane. The seat was designed for pilots weighing between 140 lb a
Answer #1

Solution :

Given that ,

mean = \mu = 146

standard deviation = \sigma = 27.2

( a ) P( 140 < x < 181 ) = P[( 141 - 146 ) / 27.2 ) < (x - \mu ) /\sigma  < ( 181 - 146 ) / 27.2) ]

= P( -0.22 < z < 1.29 )

= P( z < 1.29 ) - P( z < -0.22 )

Using z table,

= 0.9015 - 0.4129

= 0.4886

Probability = 0.4886

( b )

Given that,

n = 37

\mu\bar x = 146

\sigma\bar x = \sigma / \sqrt n = 27.2 / \sqrt 37 = 4.4717

P( 140 < T < 181 )  

= P[(140 - 146 ) / 4.4717 < (T - \mu \bar x) / \sigma \bar x < ( 181 - 146 ) / 4.4717 )]

= P( -1.34 < Z < 7.83 )

= P( Z < 7.83 ) - P(Z < -1.34 )

Using z table,  

= 1.0000 - 0.0901

= 0.9099

Probability = 0.9099  

( c )

Correct option : - ( D )